OpenStudy (liyah):

Factor X^3 + 6x^2 + 12x + 8

OpenStudy (anonymous):

(x+2)^3

jhonyy9 (jhonyy9):

x^3 + 6x(x+2) +2^3 =(x+2)(x^2 -2x +4) +6x(x+2) =(x+2)(x^2 -2x +4 +6x) = =(x+2)(x+2)^2 =(x+2)^3

jhonyy9 (jhonyy9):

- so thi is for you can understanding all work for this result